### v1

# Methodology for Fine-Grain GPU Power Visibility and Insights

Abstract
Ubiquity of AI coupled with its steep power demands make optimizing GPU power a priority as large GPU-based clusters are often employed to train and serve AI models. An important first step in optimizing GPU power consumption is high-fidelity and fine-grain power measurement of key AI computations on GPUs. To this end, we observe that as GPUs get more powerful, the resulting sub-millisecond to millisecond executions make fine-grain power analysis challenging. In this work, we first carefully identify the challenges in obtaining fine-grain GPU power profiles. To address these challenges, we devise FinGraV methodology where we employ execution time binning, careful CPU-GPU time synchronization, and power profile differentiation to collect fine-grain GPU power profiles across prominent AI computations and across spectrum of scenarios. Using FinGraV power profiles, we make several observations pertaining to GPU power variation over executions and over time, GPU sub-component power consumptions across different scenarios, and power behavior over interleaved executions of multiple computations. Equipped with these observations, we conclude with several recommendations to optimize the power for these ubiquitous accelerators.
Index Terms:
AI, fine-grain power analysis, GPU
I Introduction
A key enabler for the current AI wave has been the compute and memory horsepower availed by modern GPUs. As demand for AI sky-rockets, hyperscalers continue to deploy large-scale GPU clusters to train and serve AI models [1, 2]. This increasing demand for AI computation translates into increasing AI energy needs, which is dominated by power consumption of GPUs. As such, optimizing GPU power and energy consumption can go a long way in reducing energy consumption and capital costs of large-scale GPU clusters being deployed for AI. GPU power optimization is not only beneficial for AI but also for other domains, as the use cases for which GPUs are being utilized continue to widen. As an example, even for the HPC domain, the majority (more than 75%) of node-level power provisioning for Hewlett Packard Enterprise Frontier [3], the world’s first exascale supercomputer, is for GPUs.
An important first step in optimizing GPU power is better visibility into GPU power. Specifically, in this work, we focus on measuring fine-grain GPU power profiles along both time and space dimensions. That is, first, for fine-grain profiling along time, we focus on power measurements throughout the execution of a GPU computation (commonly referred to as a kernel in GPU parlance). Second, for fine-grain profiling along space dimension, we focus on breaking down GPU power consumption into sub-components (e.g., compute cores, HBM memory, etc.). We believe that such fine-grain visibility can go a long way in identifying avenues for and hence design of intelligent GPU power management capabilities.
Figure 1: FinGraV addresses challenges in fine-grain GPU power analysis.
However, fine-grain GPU power analysis is challenging for many reasons as depicted in Figure 1. First, as GPUs get more powerful (higher compute/memory throughputs), execution times for computations can often end up in sub-millisecond (ms) to few ms range. This makes getting fine-grain power profiles challenging even with a 1ms power sampler natively available on a GPU (C1) [4, 5, 6]. Further, a common technique of repeated kernel executions is challenging to use as kernels with such low execution times can manifest execution time variation (e.g., due to slight differences in memory allocation and hence access patterns) which must be tackled (C2). Not only do execution time variations need to be accounted for, even power variations can occur across executions, across interleaving of various kinds of computations and more (C3). Finally, using a high frequency GPU-side power sampler can either lead to repeated CPU-GPU communication for online measurements (as CPU schedules computations on the GPU today) or require careful synchronization of CPU-GPU time during post-processing of power logs obtained to identify logs that belong to the kernel execution (C4).
To tackle these challenges and provide fine-grain GPU power visibility, in this work, we propose FinGraV(Fine-Grain Visibility) methodology to collect fine-grain power profiles on the state-of-the-art AMD Instinct MI300X GPU, which is being used exclusively to serve all Llama 405B live traffic at Meta [7]. To design FinGraV, we harness GPU-side power logging and via careful synchronization of CPU-GPU time, we identify power logs of interest along with identifying its specific time in the kernel’s execution, thus getting a fine-grain profile in time. Further, we employ execution time binning strategy to tackle execution time variation. Finally, we study power behavior for a given GPU kernel under different scenarios to identify power variations that can manifest and identify specific power profiles of interest for a specific kernel.
We study the profiles that FinGraV leads to for prominent AI computations (matrix-matrix multiplication and collective communication kernels) across a spectrum of scenarios. We make several observations based on these profiles about GPU power variation over executions and over time, GPU sub-component power consumptions across different scenarios, and power behavior over interleaved executions of multiple operators. Finally, based on these observations, we conclude with a set of recommendations to optimize GPU power via both software and hardware optimizations.
Overall, we make the following contributions in this work:
  • We observe that as accelerators such as GPUs get more powerful leading to lower execution times (sub-ms to few ms), their power analysis gets increasingly challenging. To this end, we first carefully identify these challenges and their implications.
  • To address the above identified challenges, we propose FinGraV (Fine-Grain Visibility) methodology, where we employ execution time binning, careful CPU-GPU time synchronization, and power profile differentiation to collect fine-grain GPU power profiles.
  • Next, we employ FinGraV methodology to profile prominent AI computations, namely matrix-matrix multiplication (GEMM) and communication, across a spectrum of scenarios (GEMMs: compute versus memory-bound; communication: latency versus bandwidth-bound) and setups (isolated executions, interleaved executions, and more). Based on our analysis, we provide heuristics to guide fine-grain power profiling of other kernels.
  • From the above FinGraV profiles, we make several observations about GPU power variation over executions and over time, GPU sub-component power consumptions across different scenarios, and power behavior over interleaved executions of multiple computations.
  • As GPUs continually scale compute and memory throughput, power constraints will increasingly limit them. We believe FinGraV stands to provide the necessary visibility into these ubiquitous accelerators. Based on our detailed profiling, we conclude with a discussion of opportunities to optimize power for GPUs.
II Background
II-A AMD Instinct MI300X GPU and Sub-components
We focus in this work on fine-grain power analysis of the state-of-the-art AMD Instinct MI300X GPU, based on AMD CDNA 3 architecture, shown in Figure 2. As depicted, the MI300X GPU is a chiplet-based design which harnesses advanced packaging to integrate heterogeneous chiplets each specialized for a specific function. Reading the figure from the top, MI300X has accelerator complex dies (XCD) vertically stacked over I/O dies (IOD), which in turn are stacked over a passive silicon interposer. There are a total of four IODs each having two XCDs stacked for a total of eight XCDs in a single MI300X GPU [8, 9].
The IODs contain AMD Infinity Cache, a shared memory-side last-level cache (LLC) with a total capacity of 256MB. The IODs also contain the memory interface to the on-package eight stacks of high-bandwidth memory (HBM). The total HBM capacity is 192GB (24GB per HBM stack) for a combined peak memory bandwidth of 5.3TB/s [10]. The XCDs are the key computational workhorses in MI300X and each XCD in turn comprises 38 active compute units (CUs or GPU cores) for a total of 304 CUs in a single MI300X. The CUs within a single XCD share an L2 cache of 4MB (combined capacity of 32MB over eight XCDs). Computation is launched on GPUs in the form of a kernel and sub-units of a kernel (termed workgroups) are spread over available XCDs.
[…]
Figure 2: An illustration (not to scale) of AMD Instinct MI300X, the GPU used in this work. The cross-sectional view (bottom) shows the stacking.
Figure 3: Challenges in doing fine-grain GPU power analysis.
II-B ML Focus and Operators of Interest
While continued improvements in high-end GPUs have widened their use from high-performance computing (HPC) to AI and more, in this work, we focus on the usage of GPUs in the AI domain. We do so as the continued AI demand, and hence its ubiquity, has led to a concomitant increase in AI energy/power expenditure making it a great case study for power profiling and optimizations. As an example, Amazon’s training of a 200B AI model over 48 days consumed about 11.9 GWhr [11], equivalent to the average power consumption of over 1100 US households for a year [12]. That said, our proposed methodologies will equally apply to GPU kernels in other domains.
While AI workloads comprise a variety of operators, we consciously focus on two primary operators for our analysis, namely, general matrix-matrix multiplications (GEMM) kernels and communication kernels for they contribute to the majority of the AI execution time [13]. We consider both GEMM and communication kernels across a spectrum of scenarios (GEMMs: compute versus memory-bound; communication: latency versus bandwidth-bound) and setups (isolated, interleaved executions, and more) in order to study their effect on power consumption.
III Challenges to Fine-grain GPU Power Analysis
Recall that the focus of our work is to get fine-grain power profiles for GPU kernels. That is, power measurements throughout the execution of a GPU kernel (fine-grain in time dimension). Additionally, should the tooling support it, power measurement at GPU sub-component granularity (fine-grain in space dimension) such as power consumption of XCD, IOD, etc. in Figure 2. However, such fine-grain GPU power analysis in time dimension is challenging due to the following.
C1
Low sampling frequency: As compute throughput and memory bandwidth made available by high-end GPUs continue to scale, short of problem sizes scaling commensurately, GPU kernel execution times can often end up in sub-millisecond (ms) to few ms range. This is certainly true for the AI kernels we have benchmarked on MI300X GPU (Section V). This means that any CPU-side GPU power measurements at low sampling frequency, as Figure 3a depicts, can completely miss sampling power for a given kernel.
C2
CPU-GPU time synchronization: To partially overcome the above challenge, a high sampling frequency power sampler can be used. For example, on MI300X GPU, we can tap into a 1ms power logger. However, kernel scheduling events are controlled/triggered by the CPU. As such, correlating this GPU power logger (agnostic of kernel start/end) with CPU time is necessary to accurately capture the power log-of-interest (LOI) and time-of-interest (TOI) as depicted in Figure 3b.
C3
Execution time variation: Even with the above challenges addressed, with sub-ms kernel executions, a 1ms sampler will at best deliver a single power measurement requiring multiple runs to build a fine-grain power profile. However, in the sub-ms execution space, even slight variation in kernel execution time (e.g., due to slight differences in memory allocation and hence access patterns) makes correlating power measurements across runs a challenge. As depicted in Figure 3c, power measurements during three separate runs are potentially at different TOI in the kernel.
C4
Power variance across executions: Finally, note that the above execution time variation, coupled with the fact that power is continuous rather than discrete, leads to power variation manifesting as well. Specifically, as depicted in Figure 3d, we observe in this work that repeated executions of the same kernel (tagged 1/2/3 in Figure 3d) or interleaving of a kernel with other kernels (kernel 1 interleaved with kernel x and y) can lead to different power profiles. As such, identifying which power profiles to focus on is important.
IV FinGraV Methodology
Figure 4: FinGraV strategies to address challenges in fine-grain GPU power analysis.
IV-A Addressing Challenges
We discuss in this section, the broad strokes of our solution to the challenges we identified in Section III.
S1
On-GPU power logger: We harness a 1ms power logger available internally at AMD on MI300X. We discuss using the proposed FinGraV methodology in tandem with externally available power logging tools on AMD GPUs such as amd-smi [14] in Section VI.
S2
High-resolution CPU-GPU sync: As the internal power logger logs power measurements on the GPU while being agnostic of kernel start/end events as discussed in Section III, we need to employ careful syncing of CPU-GPU time to identify the power log which was taken during the execution of the kernel (LOI) and where in the kernel execution was the log taken (TOI). To do so, first, we read a GPU timestamp counter before kernel execution from the CPU side. Second, we separately benchmark the delay to read this timestamp. Finally, along with power logging, we also log this timestamp value (depicted in Figure 4b). By associating the GPU timestamp with every power log and syncing a GPU timestamp (T0) with a specific timestamp on the CPU (Tc), we can post-process the power logs to identify LOI/TOI by identifying the kernel-start/end times (also in CPU time domain) in relation to the synced CPU time Tc.
S3
Kernel execution time binning: Note that, as discussed in Section III, a single run is insufficient to create fine-grain power profiles. That said, correlating measurements across runs is challenging due to kernel execution time variation. To tackle this, we employ a simple strategy of execution time binning as depicted in Figure 4c. That is, based on our empirical experiments, we guide the user towards two heuristics for fine-grain power profiles: (1) #runs to execute and (2) margin of execution time variation to allow. By excluding outlier runs with considerable variations and fine-tuning the margin to kernel execution time (Section V-B), we lower the effects of execution time variation. Note that, outlier executions are important to study and while we focus on the common case in this work, we discuss power analysis for outlier executions in Section VI.
S4
Power profile differentiation: To tackle power variations (Figure 3d), we identify two specific power profiles of interest for a kernel based on our empirical experiments (Figure 4d). Specifically, we tag the power profile of the first kernel execution, post warm-up executions, beyond which the kernel execution time does not lower substantially (typically 3 warm-up executions from GPU idle state), as the steady-state execution (SSE) profile. This provides power visibility for the best execution time. Additionally, we tag the power profile of the kernel execution post SSE, beyond which the power does not vary substantially as the steady-state power (SSP) profile.
We observe that power can gradually rise from SSE to SSP and studying these edge points gives a potential range of power variation for a kernel in isolation, that is, when it is not perturbed by other kernels. That said, we don’t intend to convey that this range is strict or that the power of a kernel always falls in this range. As we discuss in Section V-C3, power manifested by a low-power kernel does depend heavily on kernels preceding the said kernel. As such, while we analyze both SSE/SSP profiles in this work, we also discuss the effect of interleaved kernel execution scenarios on SSE and SSP. Finally, note that, in some cases SSE and SSP profiles are the same execution/power profile. We discuss this further in Section V-D.
TABLE I: FinGraV Profiling Guidance.
Exec # Runs # LOI Binning
range margin
25-50us 400 1/5us 5%
50-200us 200 1/10us 5%
200us-1ms 200 1/10us 2%
>1ms 200 1/10us 2%
IV-B FinGraV: Steps
Bringing together the solutions identified above, we list the steps we follow in FinGraV methodology:
  1. 1.
    Time the kernel a few times (4-5) to identify the kernel execution time. Use this to lookup the FinGraV empirical guidance table (Table I) to deduce the recommended #runs, binning margin, and a guidance on #LOIs to collect for the given kernel execution time.
  2. 2.
    Add relevant CPU-side instrumentation for GPU code. This includes timing the kernel start/end, reading the GPU timestamp before kernel execution (Section IV-A), and starting/ending power logger before/after the kernel.
  3. 3.
    For the SSE profile only, per run, execute the kernel four times. For the SSP profile, do a binary search starting at four executions per run (doubling or halving #executions) to ascertain the number of executions needed to attain the SSP profile. Note that an SSP run can also get the SSE profile.
  4. 4.
    Execute specified runs with specified executions per run.
  5. 5.
    Discard all but the golden runs. Golden runs are the ones that include SSE/SSP execution times belonging to the execution time bin with the maximum number of executions within binning margin of each other (for example 5% as shown in Table I for 25-200us entries).
  6. 6.
    Synchronize CPU-GPU time to identify LOI, if available, in the collected power logs per run. Identify also the TOI (the time in kernel execution where LOI was obtained).
  7. 7.
    If #LOIs obtained thus far are less than those suggested in Table I, optionally, execute more runs (= #LOIs).
  8. 8.
    Stitch the different runs by plotting all collected LOIs and TOIs.
V FinGraV Profiles and Insights
We begin with a discussion of the AI computations under study in this work and our setup for executing them. We follow this by providing an evaluation of key tenets of FinGraV methodology and sharing some experimental profiling guidance. Finally, we discuss FinGraV profiles for AI computations, key observations from these profiles, and implications for future hardware/software based on these observations.
V-A AI Operator Space and Setup
[…]
Specifically, we cover compute-bound (CB) square (M=N=K) GEMM sizes of (8K=8192, 4K=4096, 2K=2048) and memory-bound (MB) GEMV sizes for the same matrices (i.e., M=K, N=1) for a total of six AI GEMMs. We define a kernel to be compute-bound if its algorithmic op-to-byte ratio is larger than the machine’s op-to-byte as calculated from the peak compute and memory throughput of the underlying processor (kernel is memory-bound otherwise). Additionally, for communication, we study multi-GPU collectives such as all-gather and all-reduce which are widely used in AI workloads. For collectives, we consider both latency-bound (64KB and 128KB, relevant for inference) and bandwidth-bound (512MB and 1GB, relevant for training) scenarios. Note that collective kernels, depending on associated data-transfer size, can be latency or bandwidth-bound. We classify a size as latency-bound if collective latency at/before this size does not increase commensurate to data-transfer size (kernel is bandwidth-bound otherwise).
To execute GEMMs, we harness AMD ROCm [15] rocBLAS library [16] consisting of high-performance GEMM kernels. For AI collectives, we employ AMD ROCm Communication Collectives Library (RCCL) [17], a library of standard collective communication routines for GPUs. As discussed in Section IV-B, in a given run, we can execute a kernel multiple times and collect power logs over the entire run. We use post-processing to identify executions of interest within a given run.
V-B FinGraV Methodology Evaluation
Figure 5: FinGraV methodology evaluation for (a) benefit of CPU-GPU time sync, (b) effect of kernel execution time binning, and (c) Resiliency to #runs using CB-4K-GEMM power profiles under different scenarios.
Before we present AI power profiles, we first begin with an evaluation of the key tenets of FinGraV methodology namely: (a) the benefit of CPU-GPU time sync, (b) the effect of kernel execution time binning, and (c) resiliency to #runs executed. To do so, we focus on multiple power profiles for a compute-bound (CB) 4K GEMM kernel (hence referred to as CB-4K-GEMM for simplicity) depicted in Figure 5 with and without these techniques. We depict in the figure multiple runs of a CB-4K-GEMM and multiple executions within a run (about 25). We have time for a run on the x-axis and total power profiled on the y-axis. [^1] Note that, in this work, we will focus on relative power data only (not absolute power numbers). Additionally, using vertical line markers, we separate the warmup, SSE (steady-state execution) and SSP (steady-state power) executions on the graph (Section IV-A).
Benefits of CPU-GPU Time Sync: We discussed in Section IV-A the importance of syncing CPU-GPU time while using a GPU-based power logger for events triggered by the CPU. Figure 5 shows the benefit of this time synchronization for CB-4K-GEMM by comparing the unsynchronized (red) power profile to the synchronized (blue) power profile. As shown, the synchronized profile captures the gradual rise in power as GPU moves from idle state to executing the kernel (warm-up executions, to SSE, then finally to SSP) while the unsynchronized profile misses this ramp and fails to align power changes with appropriate executions in a run. [^2] Figure 5, and subsequent profiles, depict a single SSP kernel execution, which is the last execution of a given kernel in a given run (Section IV-A).
Benefits of Kernel Execution Time Binning: We also discussed, in Section IV-A, the importance of kernel execution time binning to better tackle execution time variation. We show how this binning leads to tighter power profiles in Figure 5 where we show the profile without binning using transparent blue dots while the profile with binning is shown with filled/dark blue dots. As shown, binning leads to a tighter power profile, more tuned to the true shape of power consumed. Tighter binning margins can even further smoothen the power profile albeit at the cost of more #runs as we discuss below with the profiling guidance we offer.
TABLE II: FinGraV Profiling Insights and Recommendations for Future Hardware/Software.
# Takeaway Recommendation Section
1 Similar kernel execution times can manifest very different power profiles. Considerable smarts in power management (firmware) are necessary to be power proportional on GPUs. Section V-C1
2 Total power scales with work done and different GPU components get stressed based on computation behavior. Concurrency benefits can be reaped via stitching together kernels with complementary power profiles (given power headroom). Section V-C2
3 Compute-light kernels and compute-heavy kernels show similar compute power. Techniques to lower compute power for compute-light kernels are necessary. Section V-C2
4 Power for power-light kernels (memory-bound GEMVs and compute-light GEMMs) gets affected by kernels preceding them while power-heavy kernels are not affected. Power management mechanisms which lower the power penalty incurred by power-light kernels when they follow power-heavy kernels are necessary. Section V-C3
5 When possible, at the software-level (scheduler) or at the algorithm design, prioritizing low-power kernels in scheduling can help. Section V-C3
Resiliency to Executed #Runs: We discussed in Section IV-A that with a 1ms power logger and sub-ms kernels, we get at best a single power log in a given run. We need multiple runs to create fine-grain power profiles. While in subsequent graphs we indicate a certain #runs, we discuss here, the effect of considerably lowering the #runs on the power profile. All the CB-4K-GEMM power profiles in Figure 5 use about 200 runs but we also depict (with a dashed black line) the power profile that can be attained with 50 runs only. To get this line, we do a linear regression of degree four over the power data we get with 50 runs only. As shown, even with 50 runs, we are able to ascertain the overall power trend for CB-4K-GEMM. Consequently, while in the rest of the paper we use large #runs for smooth power profiles, fewer #runs can also be employed.
FinGraV Profiling Guidance: Finally, we show in Table I some profiling guidance largely driven by our empirical analysis for GEMM kernels. The table covers the ranges of executions we see for GEMM kernels. We observed that for smaller kernel execution times, more #runs and slightly higher kernel binning margin can be needed to get enough power LOIs to create smooth power profiles. That said, this is simply a guidance and, as we discussed above, FinGraV methodology is resilient to lowering #runs.
V-C GEMM Profiles and Insights
Figure 6: CB-8K-GEMM total and XCD power.
We discuss our GEMM power profiles (compute-bound/CB and memory-bound/MB) and learnings in this section. We start with general power trends we observe, follow that with component-level (Section II-A) comparative analysis, and finally discuss power behavior in the presence of interleaved GEMM executions. Along the way, we make several key observations and recommendations for future hardware/software to optimize GPU power (all summarized in Table II).
V-C1 Power Trends
Figure 7: CB-2K-GEMM total and XCD power.
We first begin with the general question of power trends manifested. To do so, we depict the power profile (total and XCD component) for 8K compute-bound GEMM (CB-8K-GEMM) over 200 runs with eleven executions per run in Figure 6. We make several observations here. First, the power rises for initial executions and then it drops till it reaches SSE profile. Finally, the power slightly increases till it reaches SSP profile at the eleventh execution after which it shows little variation. Also, we note that the total power is dominated by XCD power for such CB GEMMs.
We surmise that for a compute-heavy GEMM kernel such as CB-8K-GEMM, the first few executions considerably stress power, invoking the power management firmware to drop the power. Such behavior has been documented before [18]. We also observe that the best execution time, which happens for SSE execution and beyond, shows lower power than these initial executions. Finally, the SSP execution shows slightly higher power than the SSE execution.
Figure 8: Component-level comparative analysis of compute-bound GEMMs and memory-bound GEMVs.
We compare these observations to the power profile for CB-2K-GEMM (Figure 7) which is considerably less compute-heavy in comparison to CB-8K-GEMM (based on op-to-byte ratio). We observe that power starts low for initial executions before rising considerably for the SSP execution (40th execution in case of CB-2K-GEMM). Recall that all executions between SSE and SSP have similar execution times. Overall, the two key power trends we observe are (a) a sharp rise followed by a drop to a steady-state and (b) a gradual rise to a steady-state.
The above observations lead us to our first key takeaway (highlighted in Table II) that similar kernel execution times (SSE through SSP) can manifest very different power profiles. This in turn leads us to make our first recommendation that considerable smarts in power management (firmware) are necessary to ensure that the lowest power profile is exercised for a given computation.
V-C2 Component Comparative Analysis
Next, we compare component-level power across all three CB GEMMs and three MB GEMVs. The relative powers for total power and the component-level powers for XCD, IOD, and HBM is depicted in Figure 8. Note that we use SSP power profiles of these kernels to plot these graphs. That said, SSE profiles show similar trends with one exception which we highlight below. Finally, for better visibility, we also show linear regression lines for all power profiles.
We make several observations here. All CB GEMMs show considerably higher total and XCD power versus MB GEMVs. This makes sense as CB GEMMs incur heavy compute and also data movement, while MB GEMVs are data movement heavy only. [^3] As we repeatedly execute kernels, data movement is heavily biased toward on-chip data movement for our executions. Amongst CB GEMMs, CB-8K-GEMM has slightly higher total/XCD power. On the GEMV front, we see a drop in total power going from 8K-GEMV to 2K-GEMV. Unlike total/XCD power, where CB GEMMs dominate, MB-8K-GEMV does stress IOD power. Finally, as discussed above, as our data movement is biased towards on-chip data movement (repeated executions), both CB/MB GEMM kernels stress HBM power similarly (except CB-8K-GEMM, which has the highest HBM power, as its considerably large input sizes stress the on-chip caches the most).
Figure 9: Total power comparison for different interleaved GEMM and GEMV experiments.
The above observations lead us to our second key takeaway (Table II) that total power generally scales with work done (e.g., CB > MB) with different components getting stressed based on computation behavior (e.g., CB stress XCD power, MB can stress IOD power, etc.). This takeaway in turn leads to our second recommendation that, should it be possible, we should stitch together kernels with complementary power profiles at an algorithmic/software level. This, with enough power headroom, will allow us to reap the benefits of concurrent execution. An example of this for AI workloads is concurrent execution of MB attention kernel with CB fully-connected layers [19].
Figure 10: Component-level comparative analysis of the evaluated communication kernels and CB-8K-GEMM.
Another takeaway we point to is that all CB GEMMs are in the ballpark of each other when it comes to XCD power. That said, compute throughput calculated using algorithmic ops and execution times shows that CB-2K-GEMM has about half the compute utilization in comparison to CB-4K/8K-GEMM. This leads us to our third recommendation that techniques which lower the compute power for compute-light kernels (e.g., CB-2K-GEMM with comparatively lower compute utilization) are necessary. Note that, unlike XCD power, we observe that the IOD power tracks well with LLC bandwidth, underscoring the importance of focusing on the XCD component.
Finally, all of the above component-level observations also hold with SSE profiles, albeit at lower magnitude and with one exception. A compute-light kernel, such as CB-2K-GEMM, manifests power similar to MB GEMMs when it comes to its SSE profile.
V-C3 Interleaved Kernels Analysis
Next, we aim to tease out how the power profiles change when different kinds of kernels are executed in an interleaved fashion. To do so, we plot relative total powers for a given GEMM/GEMV and compare the resultant power profile to its SSP power in isolation. This is depicted in Figure 9.
Focusing on the left graph, we first compare the SSP profile of CB-8K-GEMM, a compute-heavy kernel, to its profile when it is run post 60 CB-2K-GEMMs (CB–>8K). We observe a slight rise in CB-8K-GEMM power in relation to its SSP power. Next, we compare the SSP profile of CB-2K-GEMM, a compute-light kernel, across different interleaved executions and see that there is a considerable difference in the observed power profiles. When 40 MB-4K-GEMVs are run before a single CB-2K-GEMM (MB–>2K), the power for CB-2K-GEMM is far lower than its SSP profile. In comparison, if we run CB-8K-GEMM and CB-4K-GEMM before CB-2K-GEMM (CB–>2K), its power is higher than its SSP power. This shows that the power of the compute-light kernels is affected by which kernels precede them but that of a compute-heavy kernel is not.
We see something similar for MB kernels in the graph on the right in Figure 9. MB kernels, being power-light, are affected by kernels that precede them. Comparing the SSP profile for MB-8K-GEMV, when other MB kernels (MB-4K/2K-GEMV) are interleaved before it, shows that MB-8K-GEMV consumes less power in comparison to its SSP profile (MB–>8K_gemv). On the other hand, MB-4K-GEMV shows more than its SSP power when CB kernels (CB-8K/4K-GEMM) are run before it (CB–>4K_gemv).
The above observations lead us to our fourth key takeaway that power-light kernels (MB GEMVs and compute-light GEMMs) are affected by kernels preceding them when it comes to their power consumption. This takeaway leads to two recommendations from us. First, power management mechanisms that lower the power penalty incurred by power-light kernels when they follow power-heavy kernels are necessary. This can include (but not be limited to) pro-active fine-grain voltage/frequency modulation mechanisms in power management firmware. That said, further thought is necessary in this regard on the appropriate management of power swings (high-to-low, low-to-high) as these can have non-trivial secondary effects [20]. Another recommendation is that, when possible and when we are not limited by data dependency, prioritizing the scheduling of power-light kernels over power-heavy kernels can be practiced.
V-D Communication Profiles and Insights
We next focus on power analysis for communication kernels. Recall that we profile two widely used AI communication kernels: all-gather (AG) and all-reduce (AR). Further, we exercise both latency-bound/LB (64KB/128KB) and bandwidth-bound/BB scenarios (512MB/1GB) for both kernels.
At a high level, unlike GEMM kernels, we don’t see SSE/SSP differentiation for communication kernels. As such, we focus on SSE profiles for communication kernels. Further, Figure 10 presents relative powers (total, XCD, IOD, and HBM) for all eight communication kernels. We also plot CB-8K-GEMM in the same graph for comparing GEMM and communication kernel power profiles.
We make several observations here. With regards to XCD power, CB-8K-GEMM has much higher power than communication kernels which is expected. That said, when it comes to the total power, BB communication kernels fall somewhere in the middle of LB communication and CB-GEMM kernels. This can be explained by the considerably higher IOD and HBM power incurred by BB communication kernels.
Based on the comparison of communication and computation-heavy (GEMM) kernels, our takeaway is similar to takeaway #2 in Table II that the heterogeneous power profiles manifested can be exploited for efficient concurrent execution should there be enough power headroom.
VI Discussion
FinGraV with External Power loggers: In our work, we harness an internal 1ms power logger (Section IV-A). We believe that the general FinGraV methodology we have can be applied to external power loggers available on AMD platforms such as amd-smi [14]. The key tenets of FinGraV (careful CPU-GPU time synchronization, kernel execution time binning, and power profile differentiation) are all equally relevant even with these external power loggers. All of this said, the resultant power profiles will heavily depend on the power information these external loggers report. As an example, since these loggers may report average power, any averaging done can impact the power profiles that FinGraV methodology produces. We leave further augmenting FinGraV for these tools to future work.
We believe that, in addition to FinGraV methodology, a key contribution of our work is indeed the FinGraV profiles for key AI operations of GEMMs and communication kernels. Additionally, another key contribution of our work is the recommendations we have for future software/hardware to optimize power. All of these are immensely valuable in designing power-optimized future accelerators.
Outlier Executions: We discussed in Section IV-A that we focus FinGraV profiles on the common kernel execution time and discard any outliers. While we believe that prioritizing power analysis for the common case is the right first step, understanding the power behavior of the outlier executions is also important. One way to attain FinGraV profiles for outlier executions is to employ FinGraV methodology and focus on collecting profiles for a specific outlier execution time and discarding the rest (that is changing step-5 in Section IV-B). While doable, this can be costly as more #runs can be necessary to create a large enough bin belonging to this specific outlier execution time. Another strategy could be to break down kernel execution into phases when possible and assess if this lowers variation in each phase (as compared to variation at the kernel-level). As an example, with GPU kernels, wherein each kernel launches multiple workgroups, the kernel can be artificially terminated after half the number of workgroups are completed and each half of the execution can be studied separately. We leave such investigations to future work.
VII Related Work
Power/Energy Measurement Methodologies: Prior work measured GPU power using vendor tools and interfaces [14, 21] to access native on-board telemetry readings. Tools, such as Variorum [22], harness these interfaces to provide a vendor-neutral library for power measurement across different hardware from multiple vendors. Such efforts to standardize the power measurement interface, along with exposing more telemetry and control, are critical for exascale systems as highlighted by national labs and hyperscalers [23, 24]. To sample power at a higher rate compared to vendor tools, other work physically measure power using external power meters [25]. Using these meters, prior work assessed the existing on-board power measurement and found discrepancies in the reported power, highlighting the importance of using a high-fidelity methodology for measuring and reporting power [26, 27, 28, 29, 30]. Finally, researchers built simulators and statistical models to provide fine-grain power and energy estimates [31, 32, 33, 34, 35, 36].
In this work, we focus on native on-board power measurements, instead of modeling, as the GPUs are getting more complex, pushing both compute and memory. With its fine-grain power profiles, FinGraV can be used to improve the fidelity of the power models. As discussed in Section VI, FinGraV can work with existing public power loggers (with the caveats discussed) and can complement prior power measurement methods by providing a step-by-step methodology to unlock fine-grain visibility for the evaluated kernels. This way FinGraV can be agnostic to the varying sampling rate of the power measurement interfaces and meters. As the kernel execution gets shorter, FinGraV key principles will be more critical to collect sane fine-grain power profiles.
Lang et al. [37] aimed to construct high-resolution power profiles similar to FinGraV. However, FinGraV addresses emerging trends and challenges in GPUs that prior work ignored such as execution time variation of short kernels (Section III). Unlike Lang et al., FinGraV addresses this variation using kernel execution time binning. Also, Lang et al. used repetitive CPU-GPU synchronization to address the drift between CPU and GPU clocks over time. However, the authors did not factor in the delays imposed by the CPU-GPU communication. We observed such drift and will address this challenge in future work.
Power/Energy Characterization and Optimization: Researchers used the tools and methods above to characterize the energy efficiency of critical workloads and primitives in AI and HPC running on different scales [38, 39, 40, 41, 42, 43, 44], and to study the efficiency of the latest innovations in GPUs and other accelerators [45, 46]. Prior work also investigated the impact of frequency capping, power capping, DVFS, and input data composition on energy efficiency [47, 48, 49, 50, 24, 51].
With the above characterization, researchers found optimization opportunities to improve the performance-per-watt within a single GPU and in large-scale deployments. Per-GPU optimizations focused on tuning DVFS policies during kernel execution to balance energy efficiency and performance  [52, 53, 54, 55], while other work additionally investigated the impact of tuning workload parameters [56, 57]. As for system-level optimizations, recent work identified and reduced the energy bloat during training [58, 59, 60] and inference [61] by controlling job (e.g., batch size, server instances, and model parallelism) and GPU (e.g., frequency and power cap) knobs. Other work focused on efficient power management in large-scale LLM inference by enabling power oversubscription in LLM clusters [24] or via deploying the different phases of LLM (prompt and token generation) on different machines in the cluster [62].
We harnessed FinGraV to collect fine-grain power profiles for sub-ms GEMM/V kernels and communication collectives, key AI primitives, running on AMD MI300X. These power profiles, across the sub-components of MI300X, unveiled insights related to the different power behaviors of the evaluated kernels during standalone runs and when interspersed between other kernels. This led to the key power-aware recommendations to optimize both future GPU software and hardware. Overall, using FinGraV fine-grain profiles, researchers can uncover challenges and energy bloats to better design GPU power managers to improve the energy efficiency of standalone GPUs and across large-scale clusters.
VIII Conclusion
We focus in this work on fine-grain profiling for key AI operations of matrix-matrix multiplication and communication kernels on GPUs, which are widely deployed for AI workloads. To this end, we first identified challenges in doing fine-grain GPU power profiling and proposed FinGraV methodology to address these challenges on a state-of-the-art AMD MI300X GPU. We identify several important takeaways from FinGraV profiles and conclude by making many recommendations to optimize GPU power including algorithmic/software techniques that can lower power, the necessity of focusing on lowering power for compute-light GPU kernels, and the necessity for smarts in power management infrastructure.
Acknowledgment
AMD, the AMD Arrow logo, AMD CDNA, AMD Instinct, AMD ROCm, AMD Infinity Cache, AMD Infinity Fabric, and combinations thereof are trademarks of Advanced Micro Devices, Inc. Other product names used in this publication are for identification purposes only and may be trademarks of their respective companies.